OpenStudy (anonymous):

Solve the system of equations using the elimination method. 2x + 4y = 10 4x - y = 11

OpenStudy (anonymous):

4y + 10x = 14 -4y + 2x = 10 Add the two equations 12x = 24 x = 2 From this, we obtain y = -3/2.

OpenStudy (anonymous):

2x + 4y = 10 .. (1) 4x - y = 11 .. (2) 2 X (1): 4x - 2y = 20 .. (3) (3)-(2) : -y = 9 y= -9 2x + 4(-9) = 10 x = 23

OpenStudy (anonymous):

is 23 the answer?